OpenStudy (johan14th):

Calculus 2: Arc Length help please. y=(1/12x^3)+(1/x), limit 0 to 2, however do not need to substitute in limits to equation. Thank you :).

OpenStudy (johan14th):

\[1/12x^3-1/x\]

OpenStudy (johan14th):

above statement should be + 1/x, sorry. I've only managed to take the derivative of y' = 1/4x^2 + 1/x^2

OpenStudy (phi):

arc length is \[ \int \sqrt{1+(y')^2} \ dx \] you should find \[ y'= \frac{x^2}{4} - \frac{1}{x^2} \] notice you should get a minus sign when you find the derivative of x^(-1) the next step is to square that result (I assume you can do that?) then add 1 then take the square root. (In this problem, you get a perfect square, so it works out nicely)

OpenStudy (phi):

However, if we try to evaluate the integral we will get an infinite result